OpenStudy (anonymous):

According to a magazine, the number of stores that a coffee chain operates worldwide between 1999 and 2003 is as follows where x = 0 corresponds to 1999. Year, x 0 1 2 3 4 Stores, y 2170 3629 4405 5757 7068 (a) Find an equation of the least-squares line for these data which is y=1192.4x+2221 (b) Use the result of part (a) to estimate the rate at which new stores were opened annually for the period in question.

OpenStudy (pradipgr817):

You have the answer for part a)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=1192.4x+2221

OpenStudy (pradipgr817):

x varies annually so the slope of the curve will give you the rate at which new stores opened.

OpenStudy (anonymous):

how do you find that out

OpenStudy (pradipgr817):

In a linear equation rate of change of output is given by the slope. So you have been asked to find the rate let two values be x1 abd x2 so y1=m*x1+c y2=m*x2+c now subtract y2-y1=m(x2-x1) m=(y2-y1)/(x2-x1); which is rate at which new stores open